OpenStudy (anonymous):

The area of a given hexagon is equal to the area of an equilateral triangle whose perimeter is 36 inches. Find the length of a side of the regular hexagon.

OpenStudy (ghazi):

perimeter of equilateral triangle= 3* side

OpenStudy (ghazi):

3*L=36 find L

OpenStudy (cwrw238):

area of the triangle = 1/2 * 12 * sqrt(12^2 - 6^2) area of regular hexagon = 3*sqrt3 * s^2 ------ 2 where s = length of the side equate these and solve for s

OpenStudy (cwrw238):

simplify this is 2.59808 s^2 = 6*sqrt108 s^2 = 6* sqrt108 -------- 2.59808

OpenStudy (ghazi):

@Vincent12 where are you?

OpenStudy (cwrw238):

- its logged off by the look of it - sometimes this happens i'm afraid

OpenStudy (ghazi):

you made formula of hexagon look so horrible ....he went away

OpenStudy (cwrw238):

lol

OpenStudy (ghazi):

by the way wouldn't it be easier ...if i find length of the side by perimeter and directly put that in to the area of hexagon....like by finding L =12 from above realtion?

OpenStudy (ghazi):

*relation?

OpenStudy (cwrw238):

do you mean side of hexagon = 36 / 6 = 6?

OpenStudy (ghazi):

nope after findin the length of equilateral triangle ...find the area of triangle and equate that area to the area of hexagon

OpenStudy (ghazi):

and then find the side of hexagon

OpenStudy (cwrw238):

yes - thats what i've done

OpenStudy (ghazi):

OMG!! i didn't get it either... that is really frantic

OpenStudy (cwrw238):

- don't get frantic - keep calm !!

OpenStudy (ghazi):

okay :)

OpenStudy (cwrw238):

to be honest i skipped a few steps when working out area of the triangle

OpenStudy (ghazi):

that's why it is scary...well we should leave it here....your student will come and strive to find the solution :)

OpenStudy (cwrw238):

ok